# On forbidden configurations in

### v1

@@ 1 Introduction
Let $P$ be a set of $m$ points and $\mathcal{L}$ a set of $n$ lines in the plane. We define $I(P,\mathcal{L})$ to be the set of incidences between points from $P$ and lines from $\mathcal{L}$. That is, $I(P,\mathcal{L})$ is the set of ordered pairs $(p,L)$ such that $p\in P$, $L\in\mathcal{L}$, and $p\in L$. We write $G(P,\mathcal{L})$ to denote the incidence graph for $P$ and $\mathcal{L}$. This (oriented) graph has vertex set $P\cup\mathcal{L}$ and edge set $I(P,\mathcal{L})$. More precisely, a pair $(p,L)$ with $p\in P$ and $L\in\mathcal{L}$ is an edge of $G(P,\mathcal{L})$ if and only if $p\in L$.
@@
Solymosi posed the following conjecture, which can be found in the book by Brass, Moser, and Pach [4, p. 291].
@@
This case was also considered by Suk and Tomon [19] who also posed a special variant of Conjecture 1 stating that for every fixed $k\geq 3$, every set of $n$ points and $n$ lines in the plane that do not contain a $k$-fan determines at most $o(n^{4/3})$ incidences. Here, a $k$-fan consists of $k+1$ points and $k+1$ lines such that $k$ points lie on a single line and the remaining $k$ lines connect them to the $(k+1)$st point.
Very recently, Mirzaei and Suk [10] posed the following strengthening of Conjecture 1, whose statement is also mentioned by Brass, Moser, and Pach [4, p. 291] for the configurations considered by Solymosi [17].
@@ 2 Our results
Here, we refute both Conjecture 1 and Conjecture 2 by finding sets $P_{0}$ of $2k$ points and sets $\mathcal{L}_{0}$ of $\binom{k}{2}+1$ lines such that any configuration of $n$ lattice points and $n$ lines does not contain subconfiguration isomorphic to $(P_{0},\mathcal{L}_{0})$. We call the configuration $(P_{0},\mathcal{L}_{0})$ an extended regular $k$-gon (see Section 3 for the definition). In fact, our result is more general and shows this for point sets where coordinates are algebraic numbers of bounded degrees.
@@
For every $d\in\mathbb{N}$, there is a $k=k(d)\in\mathbb{N}$ such that for all $m,n\in\mathbb{N}$, if $(P_{0},\mathcal{L}_{0})$ is the extended regular $k$-gon, then for every set $P$ of $n$ points from $V_{d}$, and for each set $\mathcal{L}$ of $m$ lines in the plane, the graph $G(P_{0},\mathcal{L}_{0})$ is not a subgraph of $G(P,\mathcal{L})$.
@@
For an integer $k\geq 3$ let $P_{0}$ be a set of $k+\binom{k}{2}$ points, where $k$ points are called black and the remaining $\binom{k}{2}$ points are white, and let $\mathcal{L}_{0}$ be a set of $2\binom{k}{2}$ lines in the plane. We call the configuration $(P_{0},\mathcal{L}_{0})$ a subdivided $k$-clique if any two black points lie on two lines from $\mathcal{L}_{0}$ that intersect in a white point, each white point lies on exactly two lines from $\mathcal{L}_{0}$, and each line from $\mathcal{L}_{0}$ contains exactly one black and one white vertex. That is, if two points from $P_{0}$ are connected by an edge if and only if they lie on a line from $\mathcal{L}_{0}$, then the resulting graph is a 1-subdivision of $K_{k}$.
@@
For an integer $k\geq 3$, there is a constant $c=c(k)$ such that any arrangement of $m$ points and $n$ lines in the plane that does not contain a subconfiguration isomorphic to a subdivided $k$-clique determines at most
$c(m^{3/4}n^{1/2}+m\log^{2}{m}+n)$
@@
For an integer $k\geq 3$, there is a constant $c=c(k)$ such that any arrangement of $n$ points and $n$ lines in the plane that does not contain a subconfiguration isomorphic to a subdivided $k$-clique determines at most $cn^{5/4}$ incidences.
@@ 3 Proof of Theorem 2
For a positive integer $n$ and a set $S$, we use $[n]$ to denote the set $\{1,\dots,n\}$ and we write $\binom{S}{n}$ for the set of all unordered $n$-tuples of distinct elements from $S$. If $p$ and $q$ are distinct points, then we use $pq$ and $\overline{pq}$ to denote the line segment and the line, respectively, determined by $p$ and $q$.
@@
See Figure 1 for an illustration. Let $R$ be the set of vertices $v_{1},\dots,v_{k}$ of a regular $k$-gon in the real projective plane. The points in $R$ determine exactly $k$ slopes and we let $t_{1},\dots,t_{k}$ be the points on the line at infinity such that every line determined by two points from $R$ contains some point $t_{i}$. We call the resulting set $Q$ of $2k$ points $v_{1},\dots,v_{k},t_{1},\dots,t_{k}$ in the real projective plane an extended regular $k$-gon. Note that $H_{k}$ is the incidence graph of $Q$ together with the set containing all lines $\overline{v_{i}v_{j}}$ and the line $L_{\infty}$ in infinity.
@@
For every integer $k\geq 3$, if $P_{0}$ is a set of points and $\mathcal{L}_{0}$ is a set of lines in the plane such that $H_{k}=(P_{0},\mathcal{L}_{0})$, then $P_{0}$ is an image of an extended regular $k$-gon via a projective transformation.
To prove this lemma, we use the following result of Jamison [7] who proved that the affine images of regular $k$-gons are the only sets of $k$ points in general position determining exactly $k$ slopes.
@@
Assume that a set $P_{0}$ of points and a set $\mathcal{L}_{0}$ of lines in the plane satisfy $H_{k}=(P_{0},\mathcal{L}_{0})$. Let $Q$ be the set containing points from $P_{0}$ that correspond to $v_{1},\dots,v_{k}$ in $H_{k}$. The set $Q$ is in general position, as $\mathcal{L}_{0}$ contains all lines determined by pairs of points from $Q$ and each of them contains exactly two points of $Q$. Let $h$ be a projective transformation that sends the line from $\mathcal{L}_{0}$ corresponding to $L_{\infty}$ in $H_{k}$ to infinity. Then, the points from $h(Q)$ are still in general position and determine exactly $k$ slopes as every line determined by points from $h(Q)$ contains one of the points $t_{i}$ and each point $t_{i}$ lies on some line determined by points from $h(Q)$.
@@
Before proving Lemma 10, we first state some auxiliary definitions. If $p$ and $q$ are two points in the real projective plane, then we denote their distance by $|pq|$. If $a$, $b$, $c$ are three distinct points on a line, then their ratio is $(a,b;c)=\xi\frac{|ac|}{|bc|}$ where $\xi=1$ if the rays from $a$ to $c$ and from $b$ to $c$ point in the same direction, and $\xi=-1$ otherwise. For four distinct collinear points $a,b,c,d$, their cross-ratio is defined as $(a,b;c,d)=\frac{(a,b;c)}{(a,b;d)}$. It is well-known that cross-ratio is invariant under projective transformations. If $a,b,c,d$ appear on the line in this order along the orientation of the line, then
@@
We show that the point $w$ has coordinates of degree at most $d^{c^{\prime}}$ for some absolute constant $c^{\prime}\geq 1$. The point $w$ is the intersection point of two lines $\overline{v_{1}v_{2}}$ and $\overline{v_{k-1},v_{k}}$. Thus, the coordinates of $w$ can be expressed in terms of the coordinates of the points $v_{1},v_{2},v_{k-1},v_{k}$ using a fixed number of additions, multiplications, and divisions. If $\alpha$ and $\beta$ are two algebraic numbers of degrees $m$ and $n$, respectively, then the degree of $\alpha+\beta$, $\alpha\cdot\beta$, and $\alpha^{-1}$ are at most $mn$, $mn$, and $m$, respectively. Since $v_{1},v_{2},v_{k-1},v_{k}\in V_{d}$, the coordinates of all points $v_{1},v_{2},v_{k-1},v_{k}$ have degree at most $d$. It follows that the degrees of the coordinates of $w$ are at most $d^{c^{\prime}}$ for some absolute constant $c^{\prime}\geq 1$.
Using the bounds on degrees of numbers obtained by additions and multiplications and the facts $f(R^{\prime})\subseteq V_{d}$ and $w\in V_{d^{c}}$, the squares of all distances between two points from $f(R^{\prime}\cup\{w\})$ are numbers of degree at most $d^{c}$ for some absolute constant $c\geq 1$. Thus, the number $(f(t_{1}),f(w);f(v_{1}),f(v_{2}))^{2}$ is has degree at most $d^{c}$. Since projective transformations preserve cross-ratios, we have
@@
Now, if $r\pi\in(0,\pi/2)$, $r$ is rational, and $\cos(r\pi)$ is a quadratic irrational, then Theorem 12 implies $r\in\left\{\frac{1}{4},\frac{1}{6},\frac{1}{5},\frac{2}{5}\right\}$.
@@
Let $x=\cos\left(\frac{2\pi}{k}\right)$ and note that $0<\frac{2\pi}{k}<\frac{\pi}{2}$ and $x>0$, since $k\geq 5$. Assume that $(1+2x)^{2}=r$ for some rational number $r$. We show that $k=6$. After rewriting, we get $04x^{2}+4x+1-r=$ 0, so $x$ is a solution to a quadratic equation with rational coefficients. In other words, $x$ is either a rational number or a real quadratic irrational.
@@
We let $p\geq 3$ be a prime number such that $(p+1)/4>d^{c}$. Let $x=\cos\left(\frac{2\pi}{p}\right)$. By Theorem 13, $x$ is an algebraic number of degree $D=\varphi(p)/2=(p-1)/2$. Suppose for contradiction that the degree of $y=(1+2x)^{2}$ is at most $(D-1)/2$. That is, there is a polynomial $P$ with rational coefficients and degree at most $(D-1)/2$ such that $P(y)=0$. Then, however, the polynomial $Q(x)=P((1+2x)^{2})$ satisfies $Q(x)=P(y)=0$, while $Q$ has rational coefficients and degree at most $12\cdot((D-1)/2)=D-$ 1. This contradicts the fact that the degree of $x$ is $D$.
@@ 4 Proof of Theorem 3
We recall the notion of the crossing number. A drawing of a simple graph $G$ in the plane is a mapping $f$ that assigns to each vertex of $G$ a distinct point of the plane, and to each edge $uv$ of $G$ a continuous arc connecting $f(u)$ and $f(v)$, without passing through the image of any other vertex. The crossing number $\textrm{cr}(G)$ of $G$ is the minimum number of edge crossing points in any drawing of $G$. By a well-known result of Ajtai at al. [1] the crossing number of $G$ with $n$ vertices and $e$ edges is at least $\Omega(e^{3}/n^{2})$, assuming that $e\geq\Omega(n)$. By putting further restrictions on the graph, one can obtain improved lower bounds for the crossing number.
@@
Let $P$ be the set of $m$ points, $\mathcal{L}$ be the set of $n$ lines, and $I=|I(P,\mathcal{L})|$. We use $P^{\prime}$ to denote the set of $n$ points dual to lines from $\mathcal{L}$ and $\mathcal{L}^{\prime}$ to denote the set of $m$ lines dual to points from $P$. Let $G$ be the graph whose vertex set is formed by the points of $P^{\prime}$, and whose edges connect the $i$th and the $(i+1)$st vertex along each line from $\mathcal{L}^{\prime}$ taken from the left where $i$ is odd. Then, $G$ has $n$ vertices and at least $(I-m)/2$ edges, and any two edges of $G$ on the same line from $\mathcal{L}^{\prime}$ are disjoint.
@@ 5 Proof of Theorem 4
$|I(P,\mathcal{L})\in|O(m^{3/4}n^{1/2}+m\log^{2}{m}+n)$
for all configurations $(P,\mathcal{L})$, where $|P|=m$ and $|\mathcal{L}|=n$, that do not contain a subdivided $k$-clique. Our approach is the same as in the proof of Theorem 3, we only use a different variant of the Crossing lemma, proved by Pach, Spencer, and Tóth [14].
A graph property $\mathcal{P}$ is monotone if whenever a graph $G$ satisfies $\mathcal{P}$, then every subgraph of $G$ also satisfies $\mathcal{P}$, and whenever graphs $G_{1}$ and $G_{2}$ satisfy $\mathcal{P}$, then their disjoint union also satisfies $\mathcal{P}$.
@@
Let $\mathcal{P}$ be a monotone graph property with $\textrm{ex}(n,\mathcal{P})\in O(n^{1+\alpha})$ for some $\alpha>0$. Then, there exists two constants $c,c^{\prime}>0$ such that for any $n$ vertex graph $G$ with property $\mathcal{P}$ and with at least $e\geq cn\log^{2}n$ edges we have
$\textrm{cr}(G)\geq c^{\prime}\frac{e^{2+1/\alpha}}{n^{1+1\alpha}}.$
We also need the following Turán-type bound by Alon, Krivelevich, and Sudakov [2].
@@
Let $H$ be a bipartite graph with maximum degree $r$ on one side. Then, there exists a constant $c=c(H)>0$ such that $ex(n,H)\leq cn^{2-1/r}$.
@@
Let $P$ be the set of $m$ points, $\mathcal{L}$ be the set of $n$ lines, and $I=|I(P,\mathcal{L})|$. Let $G$ be the graph whose vertex set is formed by the points of $P$ and whose edges connect the $i$th and the $(i+1)$st vertex along each line from $\mathcal{L}$ taken from the left where $i$ is odd. Then, $G$ has $m$ vertices and at least $(I-n)/2$ edges, and any two edges of $G$ on the same line from $\mathcal{L}$ are disjoint. We assume that $(I-n)/2\in\Omega(m\log^{2}{m})$ and $(I-n)/2\geq I/4$, as otherwise $I\in O(n)$ or $I\in O(m\log^{2}{m})$ and we are done by choosing $c$ large enough.

### v2

@@ 1 Introduction
Let $P$ be a set of $m$ points and $\mathcal{L}$ be a set of $n$ lines in the plane. We define $I(P,\mathcal{L})$ to be the set of incidences between points from $P$ and lines from $\mathcal{L}$. That is, $I(P,\mathcal{L})$ is the set of ordered pairs $(p,L)$ such that $p\in P$, $L\in\mathcal{L}$, and $p\in L$. We write $G(P,\mathcal{L})$ to denote the incidence graph for $P$ and $\mathcal{L}$. This (oriented) graph has vertex set $P\cup\mathcal{L}$ and edge set $I(P,\mathcal{L})$. More precisely, a pair $(p,L)$ with $p\in P$ and $L\in\mathcal{L}$ is an edge of $G(P,\mathcal{L})$ if and only if $p\in L$.
@@
Solymosi posed the following conjecture, which can be found in the book by Brass, Moser, and Pach [3, p. 291].
@@
This case was also considered by Suk and Tomon [19] who also posed a special variant of Conjecture 1 stating that for every fixed $k\geq 3$, every set of $n$ points and $n$ lines in the plane that do not contain a $k$-fan determines at most $o(n^{4/3})$ incidences. Here, a $k$-fan consists of $k+1$ points and $k+1$ lines such that $k$ points lie on a single line and the remaining $k$ lines connect them to the (k+1)(k+1)st point.
Very recently, Mirzaei and Suk [10] posed the following strengthening of Conjecture 1, whose statement is also mentioned by Brass, Moser, and Pach [3, p. 291] for the configurations considered by Solymosi [17].
@@ 2 Our results
Here, we refute both Conjecture 1 and Conjecture 2 by finding sets $P_{0}$ of $2k$ points and sets $\mathcal{L}_{0}$ of $\binom{k}{2}+1$ lines such that any configuration of $n$ lattice points and $n$ lines does not contain subconfiguration isomorphic to $(P_{0},\mathcal{L}_{0})$. We call the configuration $(P_{0},\mathcal{L}_{0})$ an extended regular $k$-gon (see Section 3 for the definition). In fact, our result is more general and shows this for point sets where coordinates are algebraic numbers of bounded degree.
@@
For every $d\in\mathbb{N}$, there is a $k=k(d)\in\mathbb{N}$ such that for all $m,n\in\mathbb{N}$, if $(P_{0},\mathcal{L}_{0})$ is an image of the extended regular $k$-gon via a projective transformation, then for every set $P$ of $m$ points from $V_{d}$, and for each set $\mathcal{L}$ of $n$ lines in the plane, the graph $G(P_{0},\mathcal{L}_{0})$ is not a subgraph of $G(P,\mathcal{L})$.
@@
For an integer $k\geq 3$ let $P_{0}$ be a set of $k+\binom{k}{2}$ points, where $k$ points are called black and the remaining $\binom{k}{2}$ points are white, and let $\mathcal{L}_{0}$ be a set of $2\binom{k}{2}$ lines in the plane. We call the configuration $(P_{0},\mathcal{L}_{0})$ a subdivided $k$-clique if, for any two black points, there are two lines from $\mathcal{L}_{0}$, each containing one of the black points, that intersect in a white point, each white point lies on exactly two lines from $\mathcal{L}_{0}$, and each line from $\mathcal{L}_{0}$ contains exactly one black and one white vertex. That is, if two points from $P_{0}$ are connected by an edge if and only if they lie on a line from $\mathcal{L}_{0}$, then the resulting graph is a 1-subdivision of $K_{k}$.
@@
For an integer $k\geq 3$, there are constants $c=c(k)$ and $\delta=\delta(k)$ such that any arrangement of $m$ points and $n$ lines in the plane that does not contain a subconfiguration isomorphic to a subdivided $k$-clique determines at most
$c(m^{3/4-\delta}n^{1/2-\delta}+m\log^{2}{m}+n)$
@@
For an integer $k\geq 3$, there are constants $c=c(k)$ and $\delta=\delta(k)$ such that any arrangement of $n$ points and $n$ lines in the plane that does not contain a subconfiguration isomorphic to a subdivided $k$-clique determines at most $cn^{5/4-\delta}$ incidences.
@@ 3 Proof of Theorem 2
For a positive integer $n$ and a set $S$, we use $[n]$ to denote the set $\{1,\dots,n\}$ and write $\binom{S}{n}$ for the set of all unordered $n$-tuples of distinct elements from $S$. If $p$ and $q$ are distinct points, then we use $pq$ and $\overline{pq}$ to denote the line segment and the line, respectively, determined by $p$ and $q$.
@@
See Figure 1 for an illustration. Let $R$ be the set of vertices $v_{1},\dots,v_{k}$ of a regular $k$-gon in the real projective plane. The points in $R$ determine exactly $k$ slopes, and we let $t_{1},\dots,t_{k}$ be the points on the line at infinity such that every line determined by two points from $R$ contains some point $t_{i}$. We call the resulting set $Q$ of $2k$ points $v_{1},\dots,v_{k},t_{1},\dots,t_{k}$ in the real projective plane an extended regular $k$-gon. Note that $H_{k}$ is the incidence graph of $Q$ together with the set containing all lines $\overline{v_{i}v_{j}}$ and the line $L_{\infty}$ in infinity.
@@
For every integer $k\geq 3$, if $P_{0}$ is a set of points and $\mathcal{L}_{0}$ is a set of lines in the plane such that $H_{k}=G(P_{0},\mathcal{L}_{0})$, then $P_{0}$ is an image of an extended regular $k$-gon via a projective transformation.
To prove this lemma, we use the following result of Jamison [6], who proved that the affine images of regular $k$-gons are the only sets of $k$ points in general position determining exactly $k$ slopes.
@@
Assume that a set $P_{0}$ of points and a set $\mathcal{L}_{0}$ of lines in the plane satisfy $H_{k}=(P_{0},\mathcal{L}_{0})$. Let $Q$ be the set containing points from $P_{0}$ that correspond to $v_{1},\dots,v_{k}$ in $H_{k}$. The set $Q$ is in general position, as $\mathcal{L}_{0}$ contains all lines determined by pairs of points from $Q$ and each of them contains exactly two points of $Q$. Let $h$ be a projective transformation that sends the line from $\mathcal{L}_{0}$ corresponding to $L_{\infty}$ in $H_{k}$ to infinity. Then, the points from $h(Q)$ are still in general position and determine exactly $k$ slopes as every line determined by points from $h(Q)$ contains one of the points $t_{i}$, and each point $t_{i}$ lies on some line determined by points from $h(Q)$.
@@
Before proving Lemma 10, we first state some auxiliary definitions. If $p$ and $q$ are two points in the real projective plane, then we denote their distance by $|pq|$. If $a$, $b$, $c$ are three distinct points on a line, then their ratio is $(a,b;c)=\xi\frac{|ac|}{|bc|}$ where $\xi=1$ if the rays from $a$ to $c$ and from $b$ to $c$ point in the same direction, and $\xi=-1$ otherwise. For four distinct collinear points $a,b,c,d$, their cross-ratio is defined as $(a,b;c,d)=\frac{(a,b;c)}{(a,b;d)}$. It is well-known that the cross-ratio is invariant under projective transformations. If $a,b,c,d$ appear on the line in this order along the orientation of the line, then
@@
We show that the point $w$ has coordinates of degree at most $d^{c^{\prime}}$ for some absolute constant $c^{\prime}\geq 1$. The point $w$ is the intersection point of two lines $\overline{v_{1}v_{2}}$ and $\overline{v_{k-1},v_{k}}$. Thus, the coordinates of $w$ can be expressed in terms of the coordinates of the points $v_{1},v_{2},v_{k-1},v_{k}$ using a fixed number of additions, multiplications, and divisions. If $\alpha$ and $\beta$ are two algebraic numbers of degrees $m$ and $n$, respectively, then the degree of $\alpha+\beta$, $\alpha\cdot\beta$, and $\alpha^{-1}$ are at most $mn$, mnmn, and $m$, respectively. Since $v_{1},v_{2},v_{k-1},v_{k}\in V_{d}$, the coordinates of all points $v_{1},v_{2},v_{k-1},v_{k}$ have degree at most $d$. It follows that the degrees of the coordinates of $w$ are at most $d^{c^{\prime}}$ for some absolute constant $c^{\prime}\geq 1$.
Using the bounds on degrees of numbers obtained by additions and multiplications and the facts $f(R^{\prime})\subseteq V_{d}$ and $w\in V_{d^{c}}$, the squares of all distances between two points from $f(R^{\prime}\cup\{w\})$ are numbers of degree at most $d^{c}$ for some absolute constant $c\geq 1$. Thus, the number $(f(t_{1}),f(w);f(v_{1}),f(v_{2}))^{2}$ has degree at most $d^{c}$. Since projective transformations preserve cross-ratios, we have
@@
Now, if $\pi\in(0,\pi/2), rr$ is rational, and $\cos(r\pi)$ is a quadratic irrational, then Theorem 12 implies $r\in\left\{\frac{1}{4},\frac{1}{6},\frac{1}{5},\frac{2}{5}\right\}$.
@@
Let $x=\cos\left(\frac{2\pi}{k}\right)$ and note that $0<\frac{2\pi}{k}<\frac{\pi}{2}$ and $x>0$, since $k\geq 5$. Assume that $(1+2x)^{2}=r$ for some rational number $r$. We show that $k=6$. After rewriting, we get $4x^{2}+4x+1-r=0$, so $x$ is a solution to a quadratic equation with rational coefficients. In other words, $x$ is either a rational number or a real quadratic irrational.
@@
We let $p\geq 3$ be a prime number such that $(p+1)/4>d^{c}$. Let $x=\cos\left(\frac{2\pi}{p}\right)$. By Theorem 13, $x$ is an algebraic number of degree $D=\varphi(p)/2=(p-1)/2$. Suppose, for contradiction, that the degree of $y=(1+2x)^{2}$ is at most (D1)/2(D-1)/2. That is, there is a polynomial $P$ with rational coefficients and degree at most (D1)/2(D-1)/2 such that $P(y)=0$. Then, however, the polynomial $Q(x)=P((1+2x)^{2})$ satisfies $Q(x)=P(y)=0$, while $Q$ has rational coefficients and degree at most $2\cdot((D-1)/2)=D-1$. This contradicts the fact that the degree of $x$ is $D$.
@@ 4 Proof of Theorem 3
We recall the notion of the crossing number. A drawing of a simple graph $G$ in the plane is a mapping $f$ that assigns to each vertex of $G$ a distinct point of the plane, and to each edge $uv$ of $G$ a continuous arc connecting $f(u)$ and $f(v)$, without passing through the image of any other vertex. The crossing number $\textrm{cr}(G)$ of $G$ is the minimum number of edge crossing points in any drawing of $G$. By a well-known result of Ajtai et al. [1], the crossing number of $G$ with $n$ vertices and $e$ edges is at least $\Omega(e^{3}/n^{2})$, assuming that $e\in\Omega(n)$. By putting further restrictions on the graph, one can obtain improved lower bounds for the crossing number.
@@
Let $P$ be the set of $m$ points, $\mathcal{L}$ be the set of $n$ lines, and $I=|I(P,\mathcal{L})|$. We use $P^{\prime}$ to denote the set of $n$ points dual to lines from $\mathcal{L}$ and $\mathcal{L}^{\prime}$ to denote the set of $m$ lines dual to points from $P$. Let $G$ be the graph whose vertex set is formed by the points of $P^{\prime}$, and whose edges connect the $i$th and the (i+1)(i+1)st vertex along each line from $\mathcal{L}^{\prime}$ taken from the left, where $i$ is odd. Then, $G$ has $n$ vertices and at least (Im)/2(I-m)/2 edges, and any two edges of $G$ on the same line from $\mathcal{L}^{\prime}$ are disjoint.
@@
Note that it follows from the proof that the bound on $|I(P,\mathcal{L})|$ remains valid for any configuration $(P_{0},\mathcal{L}_{0})$ such that the graph $H$ corresponding to $G(P_{0},\mathcal{L}_{0})$ in $G$ is bipartite, as $H$ is then a subgraph of $K_{t,t}$ for some $t$.
@@ 5 Proof of Theorem 4
$|I(P,\mathcal{L})|\in O(m^{3/4-\delta}n^{1/2-\delta}+m\log^{2}{m}+n)$
for some constant $\delta=\delta(k)>0$, all configurations $(P,\mathcal{L})$, where $|P|=m$ and $|\mathcal{L}|=n$, that do not contain a subdivided $k$-clique. Our approach is the same as in the proof of Theorem 3; we only use a different variant of the Crossing lemma, proved by Pach, Spencer, and Tóth [14].
A graph property $\mathcal{P}$ is monotone if whenever a graph $G$ satisfies $\mathcal{P}$, then every subgraph of $G$ also satisfies $\mathcal{P}$, and whenever graphs $G_{1}$ and $G_{2}$ satisfy $\mathcal{P}$, then their disjoint union also satisfies $\mathcal{P}$. We use $ex(n,\mathcal{P})$ to denote the maximum number of edges of an $n$-vertex graph satisfying $\mathcal{P}$. If $\mathcal{P}$ is the property that our graph does not contain a subgraph isomorphic to a fixed graph $H$, then we write $ex(n,H)$ instead of $ex(n,\mathcal{P})$.
@@
Let $\mathcal{P}$ be a monotone graph property with $ex(n,\mathcal{P})\in O(n^{1+\alpha})$ for some $\alpha>0$. Then, there exists two constants $c,c^{\prime}>0$ such that for any $n$ vertex graph $G$ with property $\mathcal{P}$ and with at least $e\geq cn\log^{2}n$ edges we have
$\textrm{cr}(G)\geq c^{\prime}\frac{e^{2+1/\alpha}}{n^{1+1/\alpha}}.$
We also use the following Turán-type bound by Janzer [7].
@@
Let $H$ be a subdivision of $K_{k}$ with $k\geq 3$. Then, there exists a constant $c=c(k)>0$ such that $ex(n,H)\leq cn^{3/2-1/(4k-6)}$.
@@
Let $P$ be the set of $m$ points, $\mathcal{L}$ be the set of $n$ lines, and $I=|I(P,\mathcal{L})|$. Let $G$ be the graph whose vertex set is formed by the points of $P$ and whose edges connect the $i$th and the (i+1)(i+1)st vertex along each line from $\mathcal{L}$ taken from the left, where $i$ is odd. Then, $G$ has $m$ vertices and at least (In)/2(I-n)/2 edges, and any two edges of $G$ on the same line from $\mathcal{L}$ are disjoint. We assume that $(I-n)/2\in\Omega(m\log^{2}{m})$ and $(I-n)/2\geq I/4$, as otherwise $I\in O(n)$ or $I\in O(m\log^{2}{m})$ and we are done by choosing $c$ large enough.